Determine whether the series is convergent or divergent by expressing Sn as a telescoping sum. Me 2 02 + 4n + 3 n = 1 convergPlease read the question and follow its directions as the way I solve this problem is important.

Please show ALL work and also give an explanation to why it is convergent or divergent, that is also important.
Answer #1

Please see the solution and if you have any doubt please comment

Thanks

Solution n²+4nt3 n-1 2 We have I dans n²+4n+3 (n+1)(n+3) n=1 n1 using bastial fraction decomposition we will wife if into telwe can write if an 2 62 +40 +3 6+0 M +1 (n+3) Therefore the partial sum tan Sn = at a ta ta t.. = ( = x + ( 73 - * + + (*) tTherefore Sn = lim E 2 n²+ 40+ no D1 = lim (-btz -63) neo value is finite 6 so series is conver onvergent 5 6 Sum of series i